# Fine-Tuning Flow Matching via Maximum Likelihood Estimation of Reconstructions

### v1

@@ 3.3 Residual MLE Fine-tuning with contraction analysis
ISS and Contraction Moreover, we employ networks with specific architectures to obtain certain desirable properties, such as Input-to-State Stability (ISS) and contraction (Figs. 1(c)-1(d)). ISS refers to the property that the model output remains stable and bounded in the presence of external inputs, thereby avoiding divergence behavior as illustrated by the blue curves in Fig. 1(c). ISS property is particularly crucial in fields such as robotic manipulation. We obviously do not want minor disturbances to be excessively amplified, causing severe jitter in movements. The precise definition of ISS is provided in Appendix A. Furthermore, the contraction property indicates that the model exhibits robustness against small disturbance noises, as shown by the green curves in Fig. 1(c) and Fig. 1(d). Each contraction region can correspond to a specific semantic meaning in the latent space. We may even achieve single-step inference without relying on distillation and without loss of accuracy by directly mapping each contraction region to its corresponding equilibrium point.
Specifically, we consider the ControlSynth Neural ODE (Mei et al., 2024) or Persidskii system (Efimov & Aleksandrov, 2021; Mei et al., 2022),
$\dot{x}(t)=A_{0}x(t)+\sum_{j=1}^{M}A_{j}f_{j}(W_{j}x(t))+g(u(t)),$ (7)
where $x_{t}:=x(t)\in\mathbb{R}^{n}$ is the robot’s state vector typically including its 3-dimensional position and orientation; the matrices $A_{\cdot}$ are with approximate dimensions; $W_{\cdot}$ are weight matrices; the visual image or state input $u_{t}:=u(t)\in U\subset\mathbb{R}^{m}$, $u\in\mathscr{L}_{\infty}^{m}$; $f_{j}=[f_{j}^{1}\dots f_{j}^{k_{j}}]^{\top}$ $\left(f_{j}:\mathbb{R}^{k_{j}}\to\mathbb{R}^{k_{j}}\right)$ and $g:U\to\mathbb{R}^{n}$ ensuring the existence of the solutions of the neural network (NN) (7) at least locally in time, and $g=[g_{1}\dots g_{n}]^{\top}$; w.l.o.g., the time $t$ is set as $t\geq 0$. The definitions of the relevant symbols and further details are provided in Appendix A. Suppose that the nonlieanr function $f$ satisfying the following conditions.
Assumption 2.
[…]
Assumption 3.
Assume that the functions $f_{j}^{i}$ are continuous and strictly increasing for any $i\in\{1,\ldots,k_{j}\}$ and $j\in\{1,\ldots,M\}$.
Assumption 2 applies to many activation functions, such as $\tanh$ and parametric ReLU. With a reordering of nonlinearities and their decomposition, there exists an index $\omega\in\{0,\ldots,M\}$ such that for all $1\leq s\leq\omega$ and $1\leq i\leq k_{s}$, $\lim\limits_{\nu\to\pm\infty}f_{s}^{i}(\nu)=\pm\infty$. Also, there exists $\zeta\in\{\omega,\ldots,M\}$ such that for all $1\leq s\leq\zeta$, $1\leq i\leq k_{s}$, we have $\lim\limits_{\nu\to\pm\infty}\int\limits_{0}^{\nu}f_{s}^{i}(r)dr=+\infty$. First we introduce ISS theorem.
Theorem 3.
Let Assumptions 2-3 be satisfied. If there exist positive semidefinite symmetric matrices $P$; positive semidefinite diagonal matrices $\{\Lambda^{i}=\text{diag}(\Lambda_{1}^{i},\ldots,\Lambda_{n}^{i})\}_{i=1}^{M}$, $\{\Xi^{s}\}_{s=0}^{M}$, $\{\Upsilon_{s,r}\}_{0\leq s<r\leq M}$; positive definite symmetric matrix $\Phi$ such that the following linear matrix inequalities hold true:
$P+\sum_{j=1}^{\zeta}\Lambda^{j}>0;\quad Q=Q^{T}\leq 0;\quad\sum_{j=1}^{M}\Upsilon_{0,j}+\sum_{s=1}^{\omega}\Xi^{s}+\sum_{s=1}^{\omega}\sum_{r=s+1}^{\omega}\Upsilon_{s,r}>0.$ (8)
@@
Proof is in Appendix B. Next, to analyse contraction, we consider another trajectory of the model (7) $\dot{y}(t)=\sum_{j=1}^{M}A_{j}f_{j}(W_{j}y(t))+g(u(t))$ with the same input but different initial conditions $y(0)\in\mathbb{R}^{n}$. Let $\xi:=y-x$. Then the corresponding error system is
$\dot{\xi}=A_{0}p_{j}(\xi)+\sum_{j=1}^{M}A_{j}p_{j}(x,\xi),$ (9)
where $p_{j}(x,\xi)=f_{j}(W_{j}(\xi+x))-f_{j}(W_{j}x)$. Note that for any fixed $x\in\mathbb{R}^{n}$, the functions $p_{j}$ in the variable $\xi\in\mathbb{R}^{n}$ satisfy the properties in Assumptions 1, 2 (with a different Lipschitz constant).
Theorem 4.
Let Assumptions 2-3 and conditions in Theorem 3 be satisfied, and in the bounded domain (determined by the ISS property), the functions $f_{j}^{i}$ are Lipschitz continuous with Lipschitz constants $L_{j}^{i}$. If there exist positive semidefinite symmetric matrices $\tilde{P}$; positive semidefinite diagonal matrices $\{\tilde{\Lambda}^{i}=\text{diag}(\tilde{\Lambda}_{1}^{i},\ldots,\tilde{\Lambda}_{n}^{i})\}_{i=1}^{M}$, $\{\tilde{\Upsilon}_{j,r}\}_{j,r=1}^{M}$, $\{\Gamma_{j}\}_{j=1}^{M}$, $\{\Omega_{j}\}_{j=1}^{M}$; positive definite symmetric matrix $\Phi$; and positive scalars $\gamma,\theta$ such that the following linear matrix inequalities hold true:
$\quad\tilde{Q}=\tilde{Q}^{T}\leq 0;\quad\Gamma_{j}-\gamma L^{j}\geq 0;\quad\Omega_{j}-\theta L^{j}\geq 0;$ (10)
@@
then system (7) (with trajectory $x$) is contracting. If we define $\tilde{V}(\zeta)=\zeta^{\top}\tilde{P}\zeta+2\sum_{j=1}^{M}\sum_{i=1}^{k_{j}}\tilde{\Lambda}^{j}_{i}\int_{0}^{W_{j}^{i}\zeta}f_{j}^{i}(s)ds$, then the contraction region of $x_{0}$ contains $\{x_{0}+\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$
Proof is in Appendix B. This theorem provides the conditions for contraction and delineates the contraction region. Theorems 3 and 4 establish the stability and contraction property for systems of the form (7) via tractable linear inequalities. In practical applications, we can strictly embed these conditions into the training process, for instance, by incorporating physics-informed loss functions. Alternatively, we may simply guide the model to select matrices $A_{\cdot}$ with predominantly negative eigenvalues, while using these conditions for theoretical guarantees and interpretability analysis. Certainly, we can also employ networks with more complex architectures (e.g., UNet and Transformer). For the contraction analysis of these more general forms, we can refer to Lohmiller & Slotine (1998) and Li et al. (2025), albeit at the expense of increased condition complexity.
4 Experiments
In this section we experimentally evaluate the benefits of fine-tuning FM. All experiment details can be found in Appendix D.
4.1 Preliminary Experiments
We first perform an experiment on unconditional CIFAR-10 generation from a Gaussian source to test the basic principles of our method. The baseline models are chosen as DDPM (Ho et al., 2020), FM with Variance Exploding (VE) path and Optimal Transport (OT) path (Lipman et al., 2023), and Stochastic Interpolants (SI) (Albergo et al., 2023). We train our fine-tuned FM and report the Fréchet inception distance (FID) in Table 1. The FID over training time is in Fig 2(a).
From Table 1, we can see that fine-tuning FM can improve the performance on FID with almost the same number of function evaluations (NFE). In Fig. 2(a), the FID of FM model has stopped decreasing by the 350,000th step. But when fine-tune the checkpoint at that step, the FID score was further reduced since MLE enables more powerful representations. More visualization results and analysis are in the appendix. These demonstrate the significance of fine-tuning FM.
Table 1: FID score and number of function evaluations (NFE) for different ODE solvers: fixed-step Euler integration with 100 and 1000 steps and adaptive integration (Hairer et al., 1993, dopri5). The adaptive solver is significantly better than the Euler solver in fewer steps. First results are from Lipman et al. (2023) and the next three from Tong et al. (2024). The two last rows report the results of our fine-tuned FM. Here ResFT-FM denotes a residual MLE fine-tuning in Section 3.3 using a simplified UNet. NFE / sample $\rightarrow$ 100 1000 Adaptive Algorithm $\downarrow$ FID FID FID NFE DDPM 7.48 274 VP-FM 7.772 4.048 4.335 525.92 OT-FM 4.640 3.822 3.655 143.00 S.I. 4.488 4.132 4.009 146.12 I-CFM 4.461 3.643 3.659 146.42 OT-CFM 4.443 3.741 3.577 133.94 FT-FM (ours) 4.451 3.620 3.496 146.42 ResFT-FM (ours) 3.553 460 .07
(a) FID of Cifar-10
(b) SR in Franka Kitchen
(c) Turn on Left Burner
Figure 2: Visualization of Experimental Results. (a) plots the FID of FM (blue curve) under different training steps. The fine-tuned The red line is the fine-tuned FID score from checkpoint at $3.5\times 10^{5}$ step under the same time consumption. (b) plots the success rate (SR) of the fine-tuned policy from checkpoint at 40004000 epoch. The red vertical bars represent the variance of SR, which gradually decreases during the training process. (c) shows the task of turning on the left burner in the Franka Kitchen environment, executed by our fine-tuned FM policy within the MuJoCo simulator.
4.2 Robotic Manipulation
In this part we investigate the performance of our method on three robot manipulation datasets which includes closed-loop 6D robot actions and gripper actions: Franka Kitchen (Gupta et al., 2020), push-T (Florence et al., 2022), and Robomimic (Mandlekar et al., 2022).
  • Push-T involves manipulating a T-shaped block to a designated target using a circular end-effector. The policy takes RGB images along with end-effector proprioception as input and produces closed-loop end-effector actions. The task is supported by a dataset of 200 human demonstrations.
  • Robomimic offers 5 different tasks with high-quality human teleoperation demonstrations. This study specifically utilizes the Transport task, containing 200 demonstrations. The policy operates on state-based inputs and outputs closed-loop control signals for robot joints and the gripper.
We learn the robot policy from expert data and evaluate it in the corresponding simulation environment in MuJoCo or Gym. We choose DDPM, DDIM and orginal FM as baseline models. The evaluation in each environment has been carried out across 500 different initial conditions. The baseline models are trained for 4500 epochs. The fine-tuned training starts from the checkpoint at 3500 epochs and runs for approximately 100 epochs to ensure a roughly consistent training duration. To achieve the desired contraction property, we incorporate a physics-informed loss term $\lambda_{\omega}\mathrm{ReLU}(\omega)$ where $\omega=\$$\sum_{i,k}\left(|A_{i}^{k,k}|-\sum_{l!=k}|A_{i}^{k,l}|\right)/\left(\sum_{l}|A_{i}^{k,l}|+\epsilon_{A}\right)$ in (7) that encourages the matrix to be negatively diagonally dominant, thereby promoting a greater number of negative eigenvalues then satisfying (8) and (10). For ablation study, instead of a structurally simplistic network (7), we utilize the standard conditional UNet from the pre-training stage and apply MLE to learn the residual components. The results are recorded in Table 2 and Fig. 2(b).
Table 2 reflects that fine-tuning can effectively enhance FM’s performance. Ablation results show that training a complex network directly with MLE does not yield satisfactory results due to the optimization challenges posed by an excessive number of free parameters. Conversely, utilizing a structurally simpler contracting network (7) results in substantially improved stability and training efficiency. In subsequent work, we will analyze the various contraction regions and attempt to use them to accelerate inference. Figure 2(b) shows the success rate (SR) after fine-tuning the FM checkpoint at 4000 epochs, where the model’s performance had reached its ceiling. We observe a significant improvement in the model’s SR after only 80 epochs. The observed variance reduction also indicates enhanced reliability of our fine-tuning strategy.
Methods (16-step) Push-Ta $\uparrow$ Push-Tb $\uparrow$ Franka Kitchen $\uparrow$ Robominic $\uparrow$
DDPM 0.8840/0.7178 0.7360/0.6100 0.9840/0.6716 0.9359/0.7168
DDIM 0.8801/0.6372 0.7490/0.6167 0.9865/0.7471 0.9334/0.7073
FM 0.9035/0.7519 0.7363/0.6218 0.9960/0.7425 0.9360/0.7289
FT-FM (ours) 0.9197/0.7885 0.7567/0.6496 0.9967/0.7822 0.9401/0.7552
ResFT-FM FMc (ours) 0.9143/0.7761 0.7452/0.6511 0.9963/0.7836 0.9385/0.7447
ResFT-FM FMd (ablation) 0.9039/0.7518 0.7371/0.6220 0.9961/0.7541 0.9362/0.7283
  • a
    asampling range: [(50, 450), (50, 450), (200, 300), (200, 300), ($\pi,\pi$)]
  • b
    bsampling range: [(50, 450), (50, 450), (100, 400), (100, 400), ($\pi,\pi$)]
  • c
    cfine-tuning using ControlSynth Neural ODE (7)
  • d
    dfine-tuning using the standard conditional UNet
Table 2: We report robot performance as (max) / (average over last checkpoint with 10 replications), each averaged across 500 environment initializations. Success rate is used for all tasks except Push-T, which uses target area coverage. For Push-T, we vary initial end-effector and T-block poses. Results are shown for the Transport task in the Robomimic benchmark.
5 Conclusion and future work
While the simulation-free optimization is simple and efficient, this paper identifies some of its inherent limitations and proposes a methodology that leverages fine-tuning to further optimize FM. Experimental results robustly demonstrate the efficacy of our fine-tuning strategy. This paper also provides several novel analytical tools for flow models. We generalizes classical numerical analysis to flow settings in which the underlying vector field is approximated via learning. We combine network designs inspired from control theory to incorporate contraction property into the FM model.
In future work, we will expand the evaluation of fine-tuned FM models across a broader range of experimental settings, including spatio-temporal data forecasting. Additionally, we plan to develop methods for efficiently incorporating LMI boundedness conditions during training. The effects of more advanced numerical schemes on flow-based inference under more quantitative forms for characterizing vector field errors besides the infinity norm will be examined both theoretically and empirically. We will also explore more fine-tuning techniques, such as LoRA, and investigate alternative structured learning systems to endow the model with desirable properties such as contraction and stability. We further propose to extend this training paradigm—combining velocity matching pre-training with reconstruction error fine-tuning—to the modeling of ODE dynamical systems.
Ethics and Reproducibility Statement
Ethics statement
This work presents a theoretical and methodological advancement in generative models, especially in robotic manipulation applications. It is evaluated on standard public datasets (CIFAR-10, Push-T, Franka Kitchen, Robomimic) and does not raise any immediate ethical concerns. We encourage responsible use of the technology.
Reproducibility Statement
We provide the following to ensure reproducibility:
  • Code: Our code is in the accompanying supplementary material.
  • Experiments: Our implementation is based on PyTorch 1.12.1 and Python 3.9. We include full training and evaluation scripts, and all hyperparameters are detailed in Section D.
  • Datasets: We use the public datasets (CIFAR-10, Push-T, Franka Kitchen, Robomimic).
  • Resources: The experiments were conducted on 8 $\times$ NVIDIA A100 GPUs.
Appendix
Appendix A Notation and Definitions
For this system (7), we have the following assumption.
Assumption A.1.
We assume that the system (7) is forward complete, i.e., for all $x_{0}\in\mathbb{R}^{n}$ and $u\in\mathcal{L}_{\infty}^{q}$, the solution $x(t,x_{0},u)$ is uniquely defined for all $t\in\mathbb{R}_{+}$.
The formal definitions of ISS and convergence are provided below.
Definition A.1.
for any $x_{0}\in\mathbb{R}^{n}$ and $u\in\mathcal{L}_{\infty}^{q}$. The model (7) is convergent if it admits a unique bounded solution for $t\in\mathbb{R}$ that is globally asymptotically stable (GAS).
Appendix B Theorem Proofs
B.1 Proof of Theorem 1
Proof.
To be clear and concise, we define and rearrange some notations. For an initial value $x_{0}$, let $x(t_{n}):=\psi_{t}(x_{0})$ and $x_{n}:=\phi_{n}(x_{0})$ represent the ground truth and the model output, respectively. Then we can derive that
$\varepsilon_{n+1}$ $=x(t_{n+1})-x_{n+1}$ (B.1)
$=x(t_{n})+\int_{t_{n}}^{t_{n+1}}u_{\tau}(x(\tau))\,d\tau-x_{n}-v_{t_{n}}(x_{n})$
$=\varepsilon_{n-1}$ $+\int_{t_{n}}^{t_{n+1}}f(\tau,x(\tau))-u_{t_{n}}(x(t_{n}))+u_{t_{n}}(x(t_{n}))-u_{t_{n}}(x_{n})+u_{t_{n}}(x_{n})-v_{t_{n}}(x_{n})\,d\tau.$
We will split the integral in expression (B.1) into three parts and estimate them separately. The first part
$\ |\int_{t_{n}}^{t_{n+1}}u_{\tau}(x(\tau))-u_{t_{n}}(x(t_{n}))\,d\tau|$ (B.2)
$=|\int_{t_{n}}^{t_{n+1}}x^{\prime}(\tau)-x^{\prime}(t_{n})\,d\tau|$
$=|\frac{1}{2}\tau_{n}^{2}x^{\prime\prime}(t_{n}+\theta(\bar{t}-t_{n}))|$
$\leq\frac{1}{2}M\tau_{n}^{2},$
where $0<\theta<1$, $\bar{t}\in(t_{n},t_{n+1})$, $M=\max_{0\leq t\leq 1}|x^{\prime\prime}(t)|$. The second and third equalities in (B.2) use the Differentiation Mean Value Theorem (MVT) and Integration MVT, respectively. Using the Lipschitz condition, the second part in (B.1) derives
The third part $|\int_{t_{n}}^{t_{n+1}}u_{t_{n}}(x_{n})-v_{t_{n}}(x_{n})d\tau|\leq\delta\tau_{n}$. Plugging these into (B.1) and taking the absolute values yield
$|\varepsilon_{n+1}|\leq(1+L_{u}\tau_{n})|\varepsilon_{n}|+\delta\tau_{n}+\frac{1}{2}M\tau_{n}^{2},$ (B.3)
or equivalently,
$|\varepsilon_{n+1}|-|\varepsilon_{n}|\leq L_{u}\tau_{n}|\varepsilon_{n}|+\delta\tau_{n}+\frac{1}{2}M\tau_{n}^{2}.$ (B.4)
Summing $n$ in (B.4) from $n=0$ to m1m-1, we have
$|\varepsilon_{m}|\leq L_{u}\Sigma_{j=0}^{m-1}\tau_{j}|\varepsilon_{j}|+\Sigma_{j=0}^{m-1}(\delta\tau_{j}+\frac{1}{2}M\tau_{j}^{2})+|\varepsilon_{0}|.$
Appling Grönwall inequality (Lemma 1) we get
$|\varepsilon_{m}|\leq\exp{(L_{u}t_{m-1})}\left(\Sigma_{j=0}^{m-1}(\delta\tau_{j}+\frac{1}{2}M\tau_{j}^{2})+|\varepsilon_{0}|\right).$
By taking $m=N$, we obtain the final result in (4). It is noted that Lemma 1 employs specialized scaling techniques designed for variable step-size schemes. On a uniform grid, i.e., $\forall\,i\in[0,N]$, $\tau_{i}=\tau_{0}$, we can have a shaper error estimate. It follows directly from (B.3) that
$|\varepsilon_{n+1}|$ $\leq(1+\tau_{0}L_{u})|\varepsilon_{n}|+R$
$=(1+\tau_{0}L_{u})^{2}|\varepsilon_{n-1}|+(1+\tau_{0}L_{u})R+R$
$\leq\cdots$
$\leq(1+\tau_{0}L_{u})^{n+1}|\varepsilon_{0}|+\left[(1+\tau_{0}L_{u})^{n}+(1+\tau_{0}L_{u})^{n-1}+\cdots+1\right]R\,$,
where $R:=\delta\tau_{0}+\frac{1}{2}M\tau_{0}^{2}$. Therefore,
$|\varepsilon_{n}|$ $\leq(1+\tau_{0}L_{u})^{n}|\varepsilon_{0}|+\left[\sum_{j=0}^{n-1}(1+\tau_{0}L_{u})^{j}\right]R$
$\leq(1+\tau_{0}L_{u})^{n}|\varepsilon_{0}|+\frac{R}{\tau_{0}L_{u}}\left[(1+\tau_{0}L_{u})^{n}-1\right].$
Considering $\exp{(n\tau_{0}L_{u})}>(1+\tau_{0}L_{u})^{n}$, we can obtain
$\varepsilon_{n}\leq\exp({L_{u}\tau_{0}n})\varepsilon_{0}+\frac{M\tau_{0}+2\delta}{2L}(\exp{(L_{u}\tau_{0}n})-1).$ (B.5)
Taking $n=N$ yields the final estimate (5).
For the sake of completeness, in the following we put the standard discrete Grönwall inequality, e.g., Liao & Zhang (2021, Lemma 3.1), and its proof.
Lemma 1.
then it holds that
Proof (Proof. ).
The principle of induction completes the proof.
B.2 Proof of Theorem 2
Proof.
Under the Assumption 1 , we have
$\log p_{1}(x|x_{0},x_{1})$ $=\log\frac{\exp\left(-\frac{1}{2}(x-x_{1})^{T}\Sigma^{-1}(x-x_{1})\right)}{\sqrt{(2\pi)^{d}|\Sigma|}}$ (B.6)
When $\sigma=\mathrm{diag}(\sigma^{1},\cdots,\sigma^{d})$ is a diagonal matrix with all positive entries, it follows from B.6 that
$\log p_{1}(x|x_{0},x_{1})=-\frac{1}{2}\sum_{i=1}^{d}\frac{(x^{i}-x_{1}^{i})^{2}}{\sigma^{i}}-\log\sqrt{(2\pi)^{d}\prod_{i=1}^{d}\sigma^{d}}$ (B.7)
In the case where $\Sigma$ is a scalar matrix, i.e., $\Sigma=\sigma I$ with $\sigma\in\mathbb{R}_{+}$, following the same procedure and discarding the irrelevant coefficient leads to the loss function (6).
B.3 Proof of Theorem 3
Similar proofs can be found in Mei et al. (2024; 2022); Efimov & Aleksandrov (2021). Here, we briefly outline the general process.
Proof (Proof of Theorem 3).
@@ B.3 Proof of Theorem 3
where the vector $W_{j}^{i}$ is the $i$-th row of the matrix $W_{j}$. It is positive definite and radially unbounded due to Finsler’s Lemma under the condition equation 8 and Assumption 2. Then, taking the derivative of $V(x)$, one has
$\dot{V}$ $=$ $\begin{bmatrix}x\\ f_{1}(W_{1}x)\\ \vdots\\ f_{M}(W_{M}x)\\ g(u)\end{bmatrix}^{\top}Q\begin{bmatrix}x\\ f_{1}(W_{1}x)\\ \vdots\\ f_{M}(W_{M}x)\\ g(u)\end{bmatrix}-x^{\top}\Xi^{0}x$
@@
$-2\sum_{s=1}^{M-1}\sum_{r=s+1}^{M}f_{s}(W_{s}x)^{\top}{W_{s}^{\top}}W_{s}\Upsilon_{s,r}W_{r}^{\top}{W_{r}}f_{r}(W_{r}x)+g(u)^{\top}\Phi g(u)$
$\leq$ $-\alpha(V)+g(u)^{\top}\Phi g(u),$
for a function $\alpha\in\mathscr{K}_{\infty}$. Under (Sontag & Wang, 1995, Theorem 1), we can verify the first condition of the ISS property due to the form of $V$, and the second relation can be recovered via $V\geq\alpha^{-1}\left(2g(u)^{\top}\Phi g(u)\right)\Rightarrow\dot{V}\leq-\frac{1}{2}\alpha(V)$. This means that the ISS property of the NN equation 7 is guaranteed, and so is the boundedness of its solution.
B.4 Proof of Theorem 4
To analyze the contraction property of (9), we need the following lemma.
Lemma 2.
Under Assumption 3, we have $p_{j}(x,\xi)^{\top}p_{j}(x,\xi)\leq\xi^{\top}W_{j}^{\top}L^{j}p_{j}(x,\xi)$ and $f_{j}(W_{j}\xi)^{\top}f_{j}(W_{j}\xi)\leq\xi^{\top}W_{j}^{\top}L^{j}f_{j}(W_{j}\xi).$

### v2

@@ A.3.3 Details and Proof of the Theorem
Assumption A.8.
Functions $f_{j}^{i}$ are continuous and strictly increasing for any $i\in\{1,\ldots,k_{j}\}$ and $j\in\{1,\ldots,M\}$.
Assumption A.7 applies to many activation functions, such as $\tanh$ and parametric ReLU. With a reordering of nonlinearities and their decomposition, there exists an index $\omega\in\{0,\ldots,M\}$ such that for all $1\leq s\leq\omega$ and $1\leq i\leq k_{s}$, $\lim\limits_{\nu\to\pm\infty}f_{s}^{i}(\nu)=\pm\infty$. Also, there exists $\zeta\in\{\omega,\ldots,M\}$ such that for all $1\leq s\leq\zeta$, $1\leq i\leq k_{s}$, we have $\lim\limits_{\nu\to\pm\infty}\int\limits_{0}^{\nu}f_{s}^{i}(r)dr=+\infty$.
Now we can introduce the ISS theorem.
Theorem A.9.
Let Assumptions A.7-A.8 be satisfied. If there exist positive semidefinite symmetric matrices $P$; positive semidefinite diagonal matrices $\{\Lambda^{i}=\text{diag}(\Lambda_{1}^{i},\ldots,\Lambda_{n}^{i})\}_{i=1}^{M}$, $\{\Xi^{s}\}_{s=0}^{M}$, $\{\Upsilon_{s,r}\}_{0\leq s<r\leq M}$; positive definite symmetric matrix $\Phi$ such that the following linear matrix inequalities hold true:
$P+\sum_{j=1}^{\zeta}\Lambda^{j}>0;\quad Q=Q^{T}\leq 0;\quad\sum_{j=1}^{M}\Upsilon_{0,j}+\sum_{s=1}^{\omega}\Xi^{s}+\sum_{s=1}^{\omega}\sum_{r=s+1}^{\omega}\Upsilon_{s,r}>0.$ (A.9)
@@
Next, to analyse contraction, we consider another trajectory of the model (A.8) $\dot{y}(t)=\sum_{j=1}^{M}A_{j}f_{j}(W_{j}y(t))+\tilde{v}(\bm{o}(t))$ with the same input but different initial conditions $y(0)\in\mathbb{R}^{n}$. Let $\xi:=y-x$. Then the corresponding error system is
$\dot{\xi}=A_{0}p_{j}(\xi)+\sum_{j=1}^{M}A_{j}p_{j}(x,\xi),$ (A.10)
where $p_{j}(x,\xi)=f_{j}(W_{j}(\xi+x))-f_{j}(W_{j}x)$. Note that for any fixed $x\in\mathbb{R}^{n}$, the functions $p_{j}$ in the variable $\xi\in\mathbb{R}^{n}$ satisfy the properties in Assumptions A.7-A.8 (with a different Lipschitz constant).
Theorem A.10.
Let Assumptions A.7-A.8 and conditions in Theorem A.9 be satisfied, and in the bounded domain (determined by the ISS property), the functions $f_{j}^{i}$ are Lipschitz continuous with Lipschitz constants $L_{j}^{i}$. If there exist positive semidefinite symmetric matrices $\tilde{P}$; positive semidefinite diagonal matrices $\{\tilde{\Lambda}^{i}=\text{diag}(\tilde{\Lambda}_{1}^{i},\ldots,\tilde{\Lambda}_{n}^{i})\}_{i=1}^{M}$, $\{\tilde{\Upsilon}_{j,r}\}_{j,r=1}^{M}$, $\{\Gamma_{j}\}_{j=1}^{M}$, $\{\Omega_{j}\}_{j=1}^{M}$; positive definite symmetric matrix $\Phi$; and positive scalars $\gamma,\theta$ such that the following linear matrix inequalities hold true:
$\quad\tilde{Q}=\tilde{Q}^{T}\leq 0;\quad\Gamma_{j}-\gamma L^{j}\geq 0;\quad\Omega_{j}-\theta L^{j}\geq 0;$ (A.11)
@@
then system (4) (with trajectory $x$) is contracting. If we define $\tilde{V}(\zeta)=\zeta^{\top}\tilde{P}\zeta+2\sum_{j=1}^{M}\sum_{i=1}^{k_{j}}\tilde{\Lambda}^{j}_{i}\int_{0}^{W_{j}^{i}\zeta}f_{j}^{i}(s)ds$, then the contraction region of $x_{0}$ contains $\{x_{0}+\xi|V(\xi)\leq\max_{\zeta\in\mathbb{R}^{d}}\tilde{V}(\zeta)\}$.
We provide the proofs of the two theorems below. First, we present Theorem A.9.
Proof of Theorem A.9.
Similar proofs can be found in Mei et al. (2024, 2022); Efimov and Aleksandrov (2021). Here, we briefly outline the general process.
@@
where the vector $W_{j}^{i}$ is the $i$-th row of the matrix $W_{j}$. It is positive definite and radially unbounded due to Finsler’s Lemma under the condition (A.9) and Assumption A.7. Then, taking the derivative of $V(x)$, one has
$\dot{V}$ $=$ $\begin{bmatrix}x\\ f_{1}(W_{1}x)\\ \vdots\\ f_{M}(W_{M}x)\\ \tilde{v}(\bm{o})\end{bmatrix}^{\top}Q\begin{bmatrix}x\\ f_{1}(W_{1}x)\\ \vdots\\ f_{M}(W_{M}x)\\ \tilde{v}(\bm{o})\end{bmatrix}-x^{\top}\Xi^{0}x$
@@
$-2\sum_{s=1}^{M-1}\sum_{r=s+1}^{M}f_{s}(W_{s}x)^{\top}{W_{s}^{\top}}W_{s}\Upsilon_{s,r}W_{r}^{\top}{W_{r}}f_{r}(W_{r}x)+\tilde{v}(\bm{o})^{\top}\Phi\tilde{v}(\bm{o})$
$\leq$ $-\alpha(V)+\tilde{v}(\bm{o})^{\top}\Phi\tilde{v}(\bm{o}),$
for a function $\alpha\in\mathscr{K}_{\infty}$. Under Sontag and Wang (1995, Theorem 1), we can verify the first condition of the ISS property due to the form of $V$, and the second relation can be recovered via $V\geq\alpha^{-1}\left(2\tilde{v}(\bm{o})^{\top}\Phi\tilde{v}(\bm{o})\right)\Rightarrow\dot{V}\leq-\frac{1}{2}\alpha(V)$. This means that the ISS property of the NN (A.8) is guaranteed, and so is the boundedness of its solution. ∎
Next, we now proceed to prove the contraction Theorem A.10. To analyze the contraction property of (A.10), first we need the following lemma.
Lemma A.11.
Under Assumption A.8, we have $p_{j}(x,\xi)^{\top}p_{j}(x,\xi)\leq\xi^{\top}W_{j}^{\top}L^{j}p_{j}(x,\xi)$ and $f_{j}(W_{j}\xi)^{\top}f_{j}(W_{j}\xi)\leq\xi^{\top}W_{j}^{\top}L^{j}f_{j}(W_{j}\xi).$